# Understanding Representation Dynamics of Diffusion Models via Low-Dimensional Modeling

### v2

@@ B.3 Weight sharing in diffusion models facilitates representation learning
In this section, we present some interesting preliminary results we found that may potentially explain why diffusion models outperform classical denoising autoencoders (DAEs) in representation learning: although both share the same denoising objective (2), diffusion models demonstrate superior feature learning capabilities largely due to their inherent weight-sharing mechanism. Specifically, by minimizing the loss across all noise levels, diffusion models enable parameter sharing and interaction among denoising subcomponents, effectively creating an implicit "ensemble" effect. This interaction enhances feature consistency and robustness across noise scales, leading to significantly improved representation quality compared to DAEs [Chen et al., 2024b], as illustrated in Figure 13.
To test this, we trained 10 DAEs, each specialized for a single noise level, alongside a DDPM-based diffusion model on CIFAR10 and CIFAR100. We compared feature quality using linear probing accuracy and feature similarity relative to the optimal features at $\sigma_{t}=0.06$ (where accuracy peaks) via sliced Wasserstein distance ($\mathrm{SWD}$) [Doan et al., 2024].
The results in Figure 13 confirm the advantage of diffusion models over DAEs. Diffusion models consistently outperform DAEs, particularly in low-noise regimes where DAEs collapse into trivial identity mappings. In contrast, diffusion models leverage weight-sharing to preserve high-quality features, ensuring smoother transitions and higher accuracy as noise increases. This advantage is further supported by the $\mathrm{SWD}$ curve, which reveals an inverse correlation between feature accuracy and feature differences. Notably, diffusion model features remain significantly closer to their optimal state across all noise levels, demonstrating superior representational capacity.
Our finding also aligns with prior results that sequentially training DAEs across multiple noise levels improves representation quality [Chandra and Sharma, 2014, Geras and Sutton, 2015, Zhang and Zhang, 2018]. Our ablation study further confirms that multi-scale training is essential for improving DAE performance on classification tasks in low-noise settings (details in Appendix C, Table 3).
Beyond the implicit feature ensembling effect, we further introduce a straightforward method that explicitly ensembles features from multiple noise levels to enhance downstream task performance. Our experiments demonstrate that this approach significantly improves robustness against label noise in classification tasks, both in pre-training and transfer learning settings. For detailed methods and results, we refer interested readers to Section B.4.
@@ B.4 Feature ensembling across timesteps improves representation robustness
Our theoretical insights in Section 3 imply that features extracted at different timesteps capture varying levels of granularity. Given the high linear separability of intermediate features, we propose a simple ensembling approach across multiple timesteps to construct a more holistic representation of the input. Specifically, in addition to the optimal timestep, we extract feature representations at four additional timesteps—two from the coarse (larger $\sigma_{t}$) and two from the fine-grained (smaller $\sigma_{t}$) end of the spectrum. We then train linear probing classifiers for each set and, during inference, apply a soft-voting ensemble by averaging the predicted logits before making a final decision.(experiment details in Appendix D)
We evaluate this ensemble method against results obtained from the best individual timestep, as well as a self-supervised method MAE [He et al., 2022], on both the pre-training dataset and a transfer learning setup. The results, reported in Table 1 and Table 2, demonstrate that ensembling significantly enhances performance for both EDM [Karras et al., 2022] and DiT [Peebles and Xie, 2023], consistently outperforming their vanilla diffusion model counterparts and often surpassing MAE. More importantly, ensembling substantially improves the robustness of diffusion models for classification under label noise.
@@
Table 1: Comparison of test performance across different methods under varying label noise levels. All compared models are publicly available and pre-trained on ImageNet-1K [Deng et al., 2009], evaluated using MiniImageNet classes. Bold font highlights the best result in each scenario.
@@
Table 2: Comparison of transfer learning performance across different methods under varying label noise levels. All compared models are publicly available and pre-trained on ImageNet-1K [Deng et al., 2009], evaluated on different downstream datasets. Bold font highlights the best result in each scenario.
@@ Appendix C Additional Experiments
Validation of $\hat{\bm{x}}_{\textit{approx}}^{\star}$ approximation in Section E.2.
Figure 14: Comparison between $\mathrm{SNR}$ calculated using the optimal model $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ and the $\mathrm{SNR}$ calculated with our approximation in Theorem 1. We generate MoLRG data and calculate $\mathrm{SNR}$ using both the corresponding optimal posterior function $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ and our approximation $\hat{\bm{x}}_{\textit{approx}}^{\star}$ from Theorem 1. Default parameters are set as $n=100$, $d=5$, $K=10$, and $\delta=0.3$. In each row, we vary one parameter while keeping the others fixed, comparing the actual and approximated $\mathrm{SNR}$.
In Theorem 2, we approximate the optimal posterior estimation function $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ using $\hat{\bm{x}}_{\textit{approx}}^{\star}$ by taking the expectation inside the softmax with respect to $\bm{x}_{t}$. To validate this approximation, we compare the $\mathrm{SNR}$ calculated from $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ and from $\hat{\bm{x}}_{\textit{approx}}^{\star}$ using the definition in 1 and (10) in Section E.2, respectively. We use a fixed dataset size of $2400$ and set the default parameters to $n=100$, $d=5$, $K=10$, and $\delta=0.3$ to generate MoLRG data. We then vary one parameter at a time while keeping the others constant, and present the computed $\mathrm{SNR}$ in Figure 14. As shown, the approximated $\mathrm{SNR}$ score consistently aligns with the actual score.
@@
In Figure 3, we show that both the classification accuracy and $\mathrm{SNR}$ exhibit a unimodal trend for the MoLRG data. To further illustrate this behavior, we provide a visualization of the posterior estimation and $\mathrm{SNR}$ at different noise scales in Figure 15. In the plot, each class is represented by a colored straight line, while deviations from these lines correspond to the $\delta$-related noise term. Initially, increasing the noise scale effectively cancels out the $\delta$-related data noise, resulting in a cleaner posterior estimation and improved probing accuracy. However, as the noise continues to increase, the class confidence rate drops, leading to an overlap between classes, which ultimately degrades the feature quality and probing performance.
@@
Throughout the empirical results presented in this paper, we consistently observe a performance gap between individual DAEs and diffusion models, especially in low-noise regions. Here, we use a DAE trained on the CIFAR10 dataset with a single noise level $\sigma=0.002$, using the NCSN++ architecture [Karras et al., 2022]. In the default setting, the DAE achieves a test accuracy of $32.3$. We then explore three methods to improve the test performance: (a) adding dropout, as noise regularization and dropout have been effective in preventing autoencoders from learning identity functions [Steck, 2020]; (b) adopting EDM-based preconditioning during training, including input/output scaling, loss weighting, etc.; and (c) multi-level noise training, in which the DAE is trained simultaneously on three noise levels $[0.002,0.012,0.102]$. Each modification is applied independently, and the results are reported in Table 3. As shown, dropout helps improve performance, but even with a dropout rate of $0.95$, the improvement is minor. EDM-based preconditioning achieves moderate improvement, while multi-level noise training yields the most promising results, demonstrating the benefit of incorporating the diffusion process in DAE training.
@@ Appendix D Experimental Details
We primarily utilize the codebase of EDM [Karras et al., 2022], which is released under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 (CC BY-NC-SA 4.0) license. We also use code from the GitHub repository accompanying [Baranchuk et al., 2022], which is licensed under the MIT License.
For the datasets, CIFAR datasets [Krizhevsky et al., 2009], ImageNet [Deng et al., 2009], Oxford 102 Flowers [Nilsback and Zisserman, 2008], and DTD [Cimpoi et al., 2014] are publicly available for academic use. The CelebA dataset [Liu et al., 2015] is released for non-commercial research purposes only under a custom license. Oxford-IIIT Pet [Parkhi et al., 2012] is available under the CC BY-SA 4.0 license. The FFHQ dataset [Karras et al., 2021] is distributed by NVIDIA under the CC BY-NC-SA 4.0 license.
@@
Most experiments are conducted on a single NVIDIA A40 GPU, except for training on subsets of images (e.g., Figure 8), which is performed using two A40 GPUs.
Experimental details for Figure 1.
@@
  • Experimental details for Figure 1(a). We train diffusion models based on the unified framework proposed by [Karras et al., 2022]. Specifically, we use the DDPM+ network, and use VP configuration for Figure 1(a). [Karras et al., 2022] has shown equivalence between VP configuration and the traditional DDPM setting, thus we call the models as DDPM* models. We train two models on CIFAR10 and CIFAR100, respectively. After training, we evaluate the learned representations via linear probing. For each noise level $\sigma(t)$, we corrupt the clean input $\bm{x}_{0}$ with Gaussian noise to obtain $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}(\bm{x}_{0}+\bm{n})$, with $\bm{n}\sim\mathcal{N}\left(\mathbf{0},\sigma_{t}^{2}\mathbf{I}\right)$ and $\sqrt{\bar{\alpha}_{t}}=1/\sqrt{\sigma_{t}^{2}+1}$. We then extract features from the decoder’s ‘16x16 block1’ layer, which consistently yields the highest classification accuracy. A logistic regression classifier is trained on these features using the training split and evaluated on the test split. We perform the linear probe for each of the following noise levels: [0.002, 0.008, 0.023, 0.060, 0.140, 0.296, 0.585, 1.088, 1.923].
@@
  • Experimental details for Figure 1(b). We exactly follow the protocol in [Baranchuk et al., 2022], using the same datasets which are subsets of CelebA [Karras et al., 2018, Liu et al., 2015] and FFHQ [Karras et al., 2021], the same training procedure, and the same segmentation networks (MLPs). The only difference is that we use a newer latent diffusion model [Rombach et al., 2022] pretrained on CelebAHQ from Hugging Face and the noise are added to the latent space. For feature extraction we concatenate the feature from the first layer of each resolution in the UNet’s decoder (after upsampling them to the same resolution as the input). We perform segmentation for each of the following noise levels:[0.010, 0.015, 0.030, 0.053, 0.098, 0.189, 0.282, 0.379, 0.480, 0.766, 1.123].
Experimental details for Figure 3 and Figure 15.
@@
For the MoLRG experiments, we train the our parameterized model (4) following the setup provided in an open-source repository [tanelp, 2022]. The model is trained on a $d=5,n=50,K=3$ and $\delta=0.2$ MoLRG dataset containing 12000 samples. Training is conducted for 200 epochs using DDPM scheduling with $T=500$, employing the Adam optimizer with a learning rate of $5\times 10^{-4}$. For $\mathrm{SNR}$ computation, we follow the definition in Section 3.3 since we have access to the ground-truth basis for the MoLRG data, i.e., $\bm{U}_{1}^{\star},\bm{U}_{2}^{\star}$, and $\bm{U}_{3}^{\star}\in\mathbb{R}^{50\times 5}$. For probing we simply train a linear probe on the feature.
@@
For both panels in Figure 3, we train our probe the same training set used for diffusion and test on five different MoLRG datasets with 9000 samples generated with five different random seeds, reporting the average accuracy and $\mathrm{SNR}$ at time steps [5, 10, 20, 40, 60, 80, 100, 120, 140, 160, 180, 240, 260, 280]. In Figure 15, we visualize the posterior estimations at time steps [5, 20, 60, 120, 260] by projecting them onto the union of the first columns of $\bm{U}_{1}^{\star},\bm{U}_{2}^{\star}$, and $\bm{U}_{3}^{\star}$ (a 3D space), then further projecting onto the 2D plane by a random $3\times 2$ matrix with orthonormal columns. The subtitles of each visualization show the corresponding $\mathrm{SNR}$ calculated as explained above.
Experimental details for Figure 5.
@@
We use pre-trained EDM models [Karras et al., 2022] for CIFAR10 and ImageNet, extracting feature representations from the best-performing layer at each timestep. For the ImageNet model, features are extracted using images from classes in TinyImageNet [Le and Yang, 2015]. Feature accuracy is evaluated via linear probing. To compute the $\mathrm{SNR}$ metric, we first normalize the extracted features by dividing each by its norm and subtracting the global mean. At each timestep, we perform class-wise SVD on the normalized features and compute $\mathrm{SNR}$ as defined in Section 3.3. Specifically, we use the top $5$ right singular vectors of each class to form $\bm{U}_{k}^{\star}$.
Experimental details for Figure 6.
@@
We use the DDPM++ network and VP configuration to train diffusion models[Karras et al., 2022] on the CIFAR10 dataset, using two network configurations: UNet-$32$ and UNet-$64$, by varying the embedding dimension of the UNet. Training dataset sizes range exponentially from $2^{8}$ to $2^{15}$. For each dataset size, both UNet-$32$ and UNet-$64$ are trained on the same subset of the training data. All models are trained with a duration of $50$M images following the EDM training setup. After training, we calculate the generalization score as described in [Zhang et al., 2023], using $10$K generated images and the full training subset to compute the score.
Experimental details for Figure 7 and Figure 8.
@@
We use the DDPM++ architecture with the EDM configuration to train a UNet-128 diffusion model [Karras et al., 2022] on CIFAR10 and CIFAR100, using 4096 image training subsets. We track the evolution of representation dynamics throughout training. For Figure 7, FID [Heusel et al., 2017] is computed using $50$K generated samples compared against the full training dataset. Classification accuracy is obtained by extracting features from the training subset and evaluating a linear classifier on the full test set. Nearest neighbors are identified by computing the smallest $\ell_{2}$ distance between each generated image and the training subset.
Experimental details for Figure 13.
@@
We train individual DAEs using the DDPM++ network and VP configuration outlined in [Karras et al., 2022] at the following noise scales:

### v3

@@ B.3 Weight sharing in diffusion models facilitates representation learning
In this section, we present some interesting preliminary results we found that may potentially explain why diffusion models outperform classical denoising autoencoders (DAEs) in representation learning: although both share the same denoising objective (2), diffusion models demonstrate superior feature learning capabilities largely due to their inherent weight-sharing mechanism. Specifically, by minimizing the loss across all noise levels, diffusion models enable parameter sharing and interaction among denoising subcomponents, effectively creating an implicit "ensemble" effect. This interaction enhances feature consistency and robustness across noise scales, leading to significantly improved representation quality compared to DAEs [chen2024deconstructing], as illustrated in Figure˜13.
To test this, we trained 10 DAEs, each specialized for a single noise level, alongside a DDPM-based diffusion model on CIFAR10 and CIFAR100. We compared feature quality using linear probing accuracy and feature similarity relative to the optimal features at $\sigma_{t}=0.06$ (where accuracy peaks) via sliced Wasserstein distance ($\mathrm{SWD}$) [doan2024assessing].
The results in Figure˜13 confirm the advantage of diffusion models over DAEs. Diffusion models consistently outperform DAEs, particularly in low-noise regimes where DAEs collapse into trivial identity mappings. In contrast, diffusion models leverage weight-sharing to preserve high-quality features, ensuring smoother transitions and higher accuracy as noise increases. This advantage is further supported by the $\mathrm{SWD}$ curve, which reveals an inverse correlation between feature accuracy and feature differences. Notably, diffusion model features remain significantly closer to their optimal state across all noise levels, demonstrating superior representational capacity.
Our finding also aligns with prior results that sequentially training DAEs across multiple noise levels improves representation quality [chandra2014adaptive, geras2014scheduled, zhang2018convolutional]. Our ablation study further confirms that multi-scale training is essential for improving DAE performance on classification tasks in low-noise settings (details in Appendix˜C, Table˜3).
Beyond the implicit feature ensembling effect, we further introduce a straightforward method that explicitly ensembles features from multiple noise levels to enhance downstream task performance. Our experiments demonstrate that this approach significantly improves robustness against label noise in classification tasks, both in pre-training and transfer learning settings. For detailed methods and results, we refer interested readers to Section˜B.4.
@@ B.4 Feature ensembling across timesteps improves representation robustness
Our theoretical insights in Section˜3 imply that features extracted at different timesteps capture varying levels of granularity. Given the high linear separability of intermediate features, we propose a simple ensembling approach across multiple timesteps to construct a more holistic representation of the input. Specifically, in addition to the optimal timestep, we extract feature representations at four additional timesteps—two from the coarse (larger $\sigma_{t}$) and two from the fine-grained (smaller $\sigma_{t}$) end of the spectrum. We then train linear probing classifiers for each set and, during inference, apply a soft-voting ensemble by averaging the predicted logits before making a final decision.(experiment details in Appendix˜D)
We evaluate this ensemble method against results obtained from the best individual timestep, as well as a self-supervised method MAE [he2022masked], on both the pre-training dataset and a transfer learning setup. The results, reported in Table˜1 and Table˜2, demonstrate that ensembling significantly enhances performance for both EDM [karras2022elucidating] and DiT [peebles2023scalable], consistently outperforming their vanilla diffusion model counterparts and often surpassing MAE. More importantly, ensembling substantially improves the robustness of diffusion models for classification under label noise.
@@
Table 1: Comparison of test performance across different methods under varying label noise levels. All compared models are publicly available and pre-trained on ImageNet-1K [deng2009imagenet], evaluated using MiniImageNet classes. Bold font highlights the best result in each scenario.
@@
Table 2: Comparison of transfer learning performance across different methods under varying label noise levels. All compared models are publicly available and pre-trained on ImageNet-1K [deng2009imagenet], evaluated on different downstream datasets. Bold font highlights the best result in each scenario.
@@ Appendix C Additional Experiments
Validation of $\hat{\bm{x}}_{\textit{approx}}^{\star}$ approximation in Section˜E.2.
Figure 14: Comparison between $\mathrm{SNR}$ calculated using the optimal model $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ and the $\mathrm{SNR}$ calculated with our approximation in Theorem˜1. We generate MoLRG data and calculate $\mathrm{SNR}$ using both the corresponding optimal posterior function $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ and our approximation $\hat{\bm{x}}_{\textit{approx}}^{\star}$ from Theorem˜1. Default parameters are set as $n=100$, $d=5$, $K=10$, and $\delta=0.3$. In each row, we vary one parameter while keeping the others fixed, comparing the actual and approximated $\mathrm{SNR}$.
In Theorem˜2, we approximate the optimal posterior estimation function $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ using $\hat{\bm{x}}_{\textit{approx}}^{\star}$ by taking the expectation inside the softmax with respect to $\bm{x}_{t}$. To validate this approximation, we compare the $\mathrm{SNR}$ calculated from $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ and from $\hat{\bm{x}}_{\textit{approx}}^{\star}$ using the definition in ˜1 and (10) in Section˜E.2, respectively. We use a fixed dataset size of $2400$ and set the default parameters to $n=100$, $d=5$, $K=10$, and $\delta=0.3$ to generate MoLRG data. We then vary one parameter at a time while keeping the others constant, and present the computed $\mathrm{SNR}$ in Figure˜14. As shown, the approximated $\mathrm{SNR}$ score consistently aligns with the actual score.
@@
In Figure˜3, we show that both the classification accuracy and $\mathrm{SNR}$ exhibit a unimodal trend for the MoLRG data. To further illustrate this behavior, we provide a visualization of the posterior estimation and $\mathrm{SNR}$ at different noise scales in Figure˜15. In the plot, each class is represented by a colored straight line, while deviations from these lines correspond to the $\delta$-related noise term. Initially, increasing the noise scale effectively cancels out the $\delta$-related data noise, resulting in a cleaner posterior estimation and improved probing accuracy. However, as the noise continues to increase, the class confidence rate drops, leading to an overlap between classes, which ultimately degrades the feature quality and probing performance.
@@
Throughout the empirical results presented in this paper, we consistently observe a performance gap between individual DAEs and diffusion models, especially in low-noise regions. Here, we use a DAE trained on the CIFAR10 dataset with a single noise level $\sigma=0.002$, using the NCSN++ architecture [karras2022elucidating]. In the default setting, the DAE achieves a test accuracy of $32.3$. We then explore three methods to improve the test performance: (a) adding dropout, as noise regularization and dropout have been effective in preventing autoencoders from learning identity functions [steck2020autoencoders]; (b) adopting EDM-based preconditioning during training, including input/output scaling, loss weighting, etc.; and (c) multi-level noise training, in which the DAE is trained simultaneously on three noise levels $[0.002,0.012,0.102]$. Each modification is applied independently, and the results are reported in Table˜3. As shown, dropout helps improve performance, but even with a dropout rate of $0.95$, the improvement is minor. EDM-based preconditioning achieves moderate improvement, while multi-level noise training yields the most promising results, demonstrating the benefit of incorporating the diffusion process in DAE training.
@@
Empirical robustness to assumption relaxation
The data assumptions in ˜1 were made to simplify the analysis and derive closed-form results. Empirically, we find that the unimodal SNR trend remains robust when these assumptions are relaxed. To demonstrate this, we train our parameterized DAE under binary MoLRG data while systematically violating each assumption. Unless otherwise stated, each experiment uses the DAE network as introduced in Section˜3.2 and a dataset contains 30,00030{,}000 samples generated with $n=50$, $d=5$, and $\delta=0.2$.
  • Overlapping class subspaces. As shown in Table˜4, we control the principal angle ($\theta$) between class subspaces and observe that overlap tends to reduce SNR across timesteps while the peak remains stable. We conjecture that overlap can be viewed as introducing additional intrinsic noise beyond the $\delta$-related term, thereby affecting the SNR value.
  • Varying subspace ranks. As shown in Table˜5, we set the class subspace dimensions to $d_{0}=10$ and $d_{1}=2$. Intuitively, the higher-rank class retains more signal and is less sensitive to noise, yielding higher and later-peaking SNR, while the low-rank class decays earlier.
  • Non-uniform mixing weights. As shown in Table˜6, we set $\pi_{0}=0.8$ and $\pi_{1}=0.2$ and observe consistently higher SNR for the majority class. We conjecture that this may stem from both the score function of the distribution and DAE training being biased toward denoising more frequent samples.
Training Setting SNR @ $\sigma_{t}$
0.008 0.023 0.060 0.140 0.296 0.585 1.088 1.923
$\theta=90^{\circ}$ (Non-overlapping) 24.88 25.15 26.95 35.84 58.27 20.84 4.05 1.57
$\theta=30^{\circ}$ (Overlapping) 16.70 16.92 18.39 25.97 38.54 22.88 14.56 12.55
Table 4: Effect of overlapping class subspaces on SNR across noise levels.
Training Setting SNR @ $\sigma_{t}$
0.002 0.008 0.023 0.060 0.140 0.296 0.585 1.088 1.923
Class 0 ($d_{0}=10$) 68.94 69.00 69.83 74.87 102.24 195.44 215.31 34.17 8.44
Class 1 ($d_{1}=2$) 7.75 7.75 7.80 8.08 8.79 6.50 1.80 0.46 0.27
Overall SNR 24.12 24.13 24.28 25.21 28.19 23.34 8.18 2.84 1.45
Table 5: Effect of subspace rank variation on SNR across noise levels.
Training Setting SNR @ $\sigma_{t}$
0.002 0.008 0.023 0.060 0.140 0.296 0.585 1.088 1.923
Class 0 ($\pi_{0}=0.8$) 52.35 52.43 52.98 56.74 75.61 134.76 72.38 13.56 4.50
Class 1 ($\pi_{1}=0.2$) 12.15 12.16 12.31 13.10 17.49 23.91 6.05 1.17 0.52
Overall SNR 38.69 38.74 39.17 41.85 55.77 90.54 35.28 7.40 2.81
Table 6: Effect of non-uniform class mixing on SNR across noise levels.
Well-separated clusters.
In our main analysis, we considered zero-mean Gaussian clusters. In this section, we show that when the clusters are well separated, the unimodal trend disappears: both classification accuracy and SNR decrease monotonically, peaking at $t=0$. In this regime, the dataset remains linearly separable throughout the denoising process, and both metrics are primarily determined by the class means. Adding noise (increasing $t$) simply blurs the class boundaries, thereby reducing separability. To verify this, we conduct an experiment on well-separated data and report the results in Table˜7, which clearly exhibit this monotonic behavior. These observations highlight the importance of overlapping clusters in giving rise to the unimodal representation dynamics.
$\sigma_{t}$ 0.030 0.053 0.098 0.189 0.282 0.379 0.480 0.588 0.704 0.830 0.989 1.492
Acc 100.0 100.0 100.0 99.8 97.7 92.4 85.7 78.5 71.8 66.0 61.6 52.6
SNR 0.540 0.539 0.535 0.524 0.513 0.502 0.484 0.465 0.436 0.406 0.370 0.297
Table 7: Classification accuracy and SNR on well-separated clusters across diffusion noise levels.
Unimodal dynamics beyond pixel space.
The unimodal dynamics are not restricted to pixel-space denoisers. In the table below, we analyze feature accuracy and SNR dynamics using the DiT-XL/2 [peebles2023scalable] model on miniImageNet, and observe in Table˜8 that the unimodal trend persists in this latent diffusion model setting as well.
$\sigma_{t}$ 0.010 0.026 0.044 0.076 0.108 0.140 0.205
Acc 73.53 74.64 75.17 75.54 75.83 75.93 76.10
SNR 0.0369 0.0372 0.0375 0.0380 0.0382 0.0384 0.0387
$\sigma_{t}$ 0.271 0.339 0.719 1.234 2.031 3.424 6.135
Acc 75.92 74.85 66.47 49.31 29.28 16.41 7.81
SNR 0.0384 0.0387 0.0350 0.0321 0.0282 0.0255 0.0211
Table 8: Feature accuracy and SNR dynamics for DiT-XL/2 on miniImageNet.
Projection-based classification analysis.
In the main paper, our linear probing experiments always use logistic regression classifiers with inner-product-based logits. Here, we re-ran the experiments in Figure˜5 using a projection-based classifier: for each class, we compute its principal subspace $\bm{V}_{1},\ldots,\bm{V}_{K}]$ from the training features, and classify each test sample by identifying the class whose subspace captures the most projection energy, computed as $\arg\max_{k}\|\bm{h}(\bm{x}_{i})\bm{V}_{k}\|^{2}$.
The results are shown in Table˜9. While the projection-based classifier yields lower overall accuracy, it reveals a more pronounced unimodal trend that aligns closely with the SNR dynamics.
Dataset Classifier Type $\sigma_{t}=0.008$ $0.023$ $0.060$ $0.140$ $0.296$ $0.585$ $1.088$ $1.923$
CIFAR10 Logistic Regression 93.72 94.94 95.20 94.11 91.36 85.25 72.95 56.21
CIFAR10 Projection 86.95 90.36 90.93 90.78 88.48 82.82 70.53 53.84
Dataset Classifier Type $\sigma_{t}=0.008$ $0.023$ $0.060$ $0.140$ $0.296$ $0.585$ $1.088$ $1.923$
TinyImageNet Logistic Regression 32.17 34.70 43.09 53.58 50.78 42.13 27.56 15.10
TinyImageNet Projection 12.89 14.38 24.10 34.78 34.88 30.02 20.99 12.64
Table 9: Logistic linear probing vs. projection-based classification accuracy on CIFAR10 (Top) and TinyImageNet (Bottom) across diffusion noise levels.
@@ Appendix D Experimental Details
We primarily utilize the codebase of EDM [karras2022elucidating], which is released under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 (CC BY-NC-SA 4.0) license. We also use code from the GitHub repository accompanying [baranchuk2021label], which is licensed under the MIT License.
For the datasets, CIFAR datasets [krizhevsky2009learning], ImageNet [deng2009imagenet], Oxford 102 Flowers [nilsback2008automated], and DTD [cimpoi14describing] are publicly available for academic use. The CelebA dataset [liu2015faceattributes] is released for non-commercial research purposes only under a custom license. Oxford-IIIT Pet [parkhi2012cats] is available under the CC BY-SA 4.0 license. The FFHQ dataset [karras2019style] is distributed by NVIDIA under the CC BY-NC-SA 4.0 license.
@@
Most experiments are conducted on a single NVIDIA A40 GPU, except for training on subsets of images (e.g., Figure˜8), which is performed using two A40 GPUs.
Experimental details for Figure˜1.
@@
  • Experimental details for Figure˜1(a). We train diffusion models based on the unified framework proposed by [karras2022elucidating]. Specifically, we use the DDPM+ network, and use VP configuration for Figure˜1(a). [karras2022elucidating] has shown equivalence between VP configuration and the traditional DDPM setting, thus we call the models as DDPM* models. We train two models on CIFAR10 and CIFAR100, respectively. After training, we evaluate the learned representations via linear probing. For each noise level $\sigma(t)$, we corrupt the clean input $\bm{x}_{0}$ with Gaussian noise to obtain $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}(\bm{x}_{0}+\bm{n})$, with $\bm{n}\sim\mathcal{N}\left(\mathbf{0},\sigma_{t}^{2}\mathbf{I}\right)$ and $\sqrt{\bar{\alpha}_{t}}=1/\sqrt{\sigma_{t}^{2}+1}$. We then extract features from the decoder’s ‘16x16 block1’ layer, which consistently yields the highest classification accuracy. A logistic regression classifier is trained on these features using the training split and evaluated on the test split. We perform the linear probe for each of the following noise levels: [0.002, 0.008, 0.023, 0.060, 0.140, 0.296, 0.585, 1.088, 1.923].
@@
  • Experimental details for Figure˜1(b). We exactly follow the protocol in [baranchuk2021label], using the same datasets which are subsets of CelebA [karras2018progressive, liu2015faceattributes] and FFHQ [karras2019style], the same training procedure, and the same segmentation networks (MLPs). The only difference is that we use a newer latent diffusion model [rombach2022high] pretrained on CelebAHQ from Hugging Face and the noise are added to the latent space. For feature extraction we concatenate the feature from the first layer of each resolution in the UNet’s decoder (after upsampling them to the same resolution as the input). We perform segmentation for each of the following noise levels:[0.010, 0.015, 0.030, 0.053, 0.098, 0.189, 0.282, 0.379, 0.480, 0.766, 1.123].
Experimental details for Figure˜3 and Figure˜15.
@@
For the MoLRG experiments, we train the our parameterized model (4) following the setup provided in an open-source repository [tiny_diffusion_repo]. The model is trained on a $d=5,n=50,K=3$ and $\delta=0.2$ MoLRG dataset containing 12000 samples. Training is conducted for 200 epochs using DDPM scheduling with $T=500$, employing the Adam optimizer with a learning rate of $5\times 10^{-4}$. For $\mathrm{SNR}$ computation, we follow the definition in Section˜3.3 since we have access to the ground-truth basis for the MoLRG data, i.e., $\bm{U}_{1}^{\star},\bm{U}_{2}^{\star}$, and $\bm{U}_{3}^{\star}\in\mathbb{R}^{50\times 5}$. For probing we simply train a linear probe on the feature.
@@
For both panels in Figure˜3, we train our probe the same training set used for diffusion and test on five different MoLRG datasets with 9000 samples generated with five different random seeds, reporting the average accuracy and $\mathrm{SNR}$ at time steps [5, 10, 20, 40, 60, 80, 100, 120, 140, 160, 180, 240, 260, 280]. In Figure˜15, we visualize the posterior estimations at time steps [5, 20, 60, 120, 260] by projecting them onto the union of the first columns of $\bm{U}_{1}^{\star},\bm{U}_{2}^{\star}$, and $\bm{U}_{3}^{\star}$ (a 3D space), then further projecting onto the 2D plane by a random $3\times 2$ matrix with orthonormal columns. The subtitles of each visualization show the corresponding $\mathrm{SNR}$ calculated as explained above.
Experimental details for Figure˜5.
@@
We use pre-trained EDM models [karras2022elucidating] for CIFAR10 and ImageNet, extracting feature representations from the best-performing layer at each timestep. For the ImageNet model, features are extracted using images from classes in TinyImageNet [le2015tiny]. Feature accuracy is evaluated via linear probing. To compute the $\mathrm{SNR}$ metric, we first normalize the extracted features by dividing each by its norm and subtracting the global mean. At each timestep, we perform class-wise SVD on the normalized features and compute $\mathrm{SNR}$ as defined in Section˜3.3. Specifically, we use the top $5$ right singular vectors of each class to form $\bm{U}_{k}^{\star}$.
Experimental details for Figure˜6.
@@
We use the DDPM++ network and VP configuration to train diffusion models[karras2022elucidating] on the CIFAR10 dataset, using two network configurations: UNet-$32$ and UNet-$64$, by varying the embedding dimension of the UNet. Training dataset sizes range exponentially from $2^{8}$ to $2^{15}$. For each dataset size, both UNet-$32$ and UNet-$64$ are trained on the same subset of the training data. All models are trained with a duration of 5050M images following the EDM training setup. After training, we calculate the generalization score as described in [zhang2024emergence], using 1010K generated images and the full training subset to compute the score.
Experimental details for Figure˜7 and Figure˜8.
@@
We use the DDPM++ architecture with the EDM configuration to train a UNet-128 diffusion model [karras2022elucidating] on CIFAR10 and CIFAR100, using 4096 image training subsets. We track the evolution of representation dynamics throughout training. For Figure˜7, FID [heusel2017gans] is computed using 5050K generated samples compared against the full training dataset. Classification accuracy is obtained by extracting features from the training subset and evaluating a linear classifier on the full test set. Nearest neighbors are identified by computing the smallest $\ell_{2}$ distance between each generated image and the training subset.
Experimental details for Figure˜13.
@@
We train individual DAEs using the DDPM++ network and VP configuration outlined in [karras2022elucidating] at the following noise scales: